# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ 6.4.2. Effects on Gender and Benevolence Jointly
[…]
Here, we show the breakdown of character gender with different suggestions for participants who marked straight women as more vs less competent than men. The breakdowns are generally quite similar, but we do note how there are more fem-coded doctors and detectives (when suggested) in stories written by participants who marked straight women as more competent.
Beyond gender on it’s own, we also consider the benevolence of the presidents (9(b)). For each suggestions condition and each potential set of attributes that could be given to the president character, we consider whether adding that suggestion type changes the proportion of presidents that have that set of attributes. We observe that overall, the model was not successful in convincing participants to make threatening president characters, with the rate of threatening president characters (regardless of gender) being quite low regardless of the presence or type of suggestions with $22/412$ stories containing a threatening president (with $8/162$ stories written with threatening suggestions of either gender having a threatening president).
We do see that when provided with benevolent or threatening fem-coded suggestions, participants wrote significantly more benevolent fem-coded presidents (benevolent fem-coded suggestions: $t(154)=3.068$, $p_{\text{FDR}}\approx 0.0134$; threatening fem-coded suggestions: $t(152)=2.526$, $p_{\text{FDR}}\approx 0.0465$).
These results weakly support H1. Joint gender and benevolence suggestions generally did not change the distribution of president characters’ attributes when comparing to stories written without suggestions. While fem-coded suggestions successfully increased the frequency of fem-coded president characters, participants did not accept these fem-coded characters being anti-stereotypically threatening, leading to an increase in benevolent fem-coded presidents when writing with either benevolent or threatening fem-coded suggestions when comparing to stories written with no suggestions.
@@
    2 (Limited working) $t(17403)=-8.97$ $0.0000$
    3 (Professional working) $t(17673)=-8.228$ $0.0000$
    4 (Full professional) $t(18053)=-10.37$ $0.0000$
    5 (Primary fluency/bilingual) $t(25097)=-8.451$ $0.0000$
@@
    5 (Primary fluency/bilingual) $t(25568)=3.258$ $0.0066$
Figure 11. AI reliance broken down by self-reported English proficiency. a) We break down the set of words in stories written by participants based on their source (e.g., the participant chose a model suggestion by pressing a suggestion button). b) We compare the proportion of model suggested words to human written or edited words in pairs of proficiency levels.
[…]
Breakdown of reliance by English proficiency. We see that participants in the “elementary proficiency” category are much more likely to accept model suggestions (via pressing a suggestion button or by typing out a model suggestions exactly).
6.5. Additional Effects: Time to Decide and Individual Differences
Beyond the primary hypotheses, here we consider the effects of suggestion type on the time to make individual decisions (expanded on in subsection B.1), effects of participant’s views and stereotypes on story attributes (expanded on with other individual differences in subsection B.3), and the effect of English proficiency on participants’ reliance on model suggestions. In the appendix, we further consider the effects of suggestions on overall story length and time to write (subsection B.2, the distribution and correlation of participant stereotypes (subsection B.4), the effect of suggestions on toxicity, sentiment, and characters’ agency in the narratives (subsection B.5).
6.5.1. Effects of Suggestion Type on Time to Make Decisions
In the Detectives scenario, we find that participants take significantly less time to decide whether accept suggestions of the partner being masc-coded when the partner is untrustworthy as opposed to trustworthy ($t(306)=-3.544$, $p_{\text{FDR}}\approx 0.0030$). Further, when the partner is untrustworthy, it takes significantly less time for participants to make decisions about masc-coded suggestions than fem-coded ones ($t(310)=-3.48$, $p_{\text{FDR}}\approx 0.0036$). This could mean that masc-coded untrustworthy detective partners are the least unexpected group. This finding is in line with what we saw about the rate of masc-coded untrustworthy detectives in stories written without suggestions and is also in line with Cao et al. (2022b)’s findings about gender-trustworthiness stereotypes in people.
@@ 6.5.2. Effect of Participants’ Views on Gender
Here we consider how participants’ views on gender affect the stories they write with and without suggestions (H4). In subsection B.3 in the appendix, we similarly consider the effects of participants’ gender identity.
For the effect of beliefs about gender, we consider the “comptence” axis in scenarios where the participant writes about a doctor, president, or detective, hypothesizing that participants who believe that women are more competent than men will write more stories about fem-coded doctors, presidents, and detectives without suggestions and be more likely to accept model suggestions of these characters being fem-coded.
We focus this analysis on the beliefs about straight men and women as participants likely defaulted to these characters being straight. We remove the $42\%$ of participants who scored straight men and women’s competence within 10 points of each other (on a 0-100 scale) and then split participants into those who marked straight women as more competent ($53\%$ of the included participants) vs less competent ($47\%$ of the included participants) than straight men. We plot the breakdown of character gender for these two groups in each scenario in Figure 10.
For all scenarios, when we compare the gendered competence groups, we see no significant difference in the fraction of masc-coded or fem-coded characters written without suggestions or with masc-coded suggestions. We do, however, observe a higher rate of fem-coded characters when suggested in the doctors and Detectives scenarios. This trend is significant for the Detectives scenario ($t(97)=2.742$, $p_{\text{FDR}}\approx 0.0304$) and marginally significant for the Doctor scenario ($t(103)=2.279$, $p_{\text{FDR}}\approx 0.0802$). This provides some evidence that participants are more willing to accept anti-stereotypical suggestions when they (or their close friends) hold anti-stereotypical beliefs.
@@ 7. Discussion, Limitations, and Implications
In this work, we examined the effect of biased predictive text suggestions on human-AI co-written text. While previous work has found that model suggestions can influence writing in multiple directions (e.g., positive vs negative sentiment, arguing social media is good vs bad for society, etc), we find evidence that, in certain writing scenarios, people are not equally influenced by pro-stereotypical and anti-stereotypical suggestions. We see that while anti-stereotypical suggestions may increase the proportion of anti-stereotypical writing, this is often not consistent enough to offset pro-stereotypical biases. The contrast between our findings and findings from prior work may be due to the stickiness of stereotypes that people hold, which are often less malleable than overt beliefs about the world (Dasgupta and Greenwald, 2001; Lai et al., 2014; Payne and Gawronski, 2010; Howe and Krosnick, 2017; Lewandowsky et al., 2012); it may also be because single word suggestions may have a smaller measured effect on people’s use of co-writing systems than full sentence or paragraph suggestions, as sentence-level suggestions have less influence on users’ views than paragraph-level suggestions (Dhillon et al., 2024).
As a result, our findings show that we should not expect co-writing with a “perfectly debiased” predictive text system to lead to “perfectly unbiased” stories—in fact, we often do not even observe parity in situations where the language model is completely skewed in an anti-stereotypical direction. For example, we find that if people are shown suggestions that describe a detective character as masc-coded or fem-coded at an equal rate, we would expect about $25.5\%$ of detectives with a specified gender to be fem-coded.
Our study has several limitations. The most significant is that participants were being asked to write a story that was not “theirs”: they may not have had a clear picture in their minds as they started writing, potentially leading to over-estimated influence. Moreover, although our study focused on creative writing scenarios, predictive text—especially on phones—is often used for other purposes. For writing that is more grounded in the real world, a predictive text system may affect how one describes an appointment with a doctor that the author had an actual appointment with (e.g., their disposition, their agency, etc.) but is unlikely to influence how one describes the doctor’s gender. Finally, our study was limited to writing in English and focused on writing tasks surrounding biases held by people in the United States (Cao et al., 2022b) and did not study cross-culture or cross-linguistic effects.
@@
Based on our findings, we caution that choosing a “debiased” language model when designing a co-writing system may not reliably yield “unbiased” text. If one is designing a system to help people write less biased text, such as system may need to provide more explicit direction to its users to achieve that outcome. Any increase in pro-stereotypical co-written stories has the potential to impact the people who read (and perhaps even write) them by forming or solidifying stereotypical associations. This is especially concerning for children who are still forming their beliefs about the world (Peterson and Lach, 1990) and for non-native speakers who generally accept more model suggestions (Buschek et al., 2021). Finally, if stories written by people using predictive text are used to train future language models, these co-written stories could lead even an “unbiased” predictive text system to become increasingly biased over time as people over-reject increasingly rare anti-stereotypical suggestions, a challenge that would not be solved with watermarking because the resulting text is human-written.
@@ A.1. Scenario: Doctor
In this scenario, the story prefix describes the speaker visiting the doctor. The model suggests that the doctor is of a particular gender and is either “confident” or “unconfident” (see examples in Table 5). According to Cao et al. (2022b), American annotators view men as comparatively more confident than women.
@@ A.1.1. Effects on Gender Alone
First, we consider how suggestions affect how participants specify a doctor character’s gender alone (Table 3). At the level of stories, we hypothesized that participants would default to making the doctor masc-coded, leading the no suggestions condition to be similar to the masc-coded suggestions condition. However, we do see marginally more masc-coded doctors when participants are given masc-coded suggestions than no suggestions ($t(230)=2.467$, $p_{\text{FDR}}\approx 0.0521$). We believe this may be due to an overall lower rate of specifying the doctor’s gender in the no suggestions condition. When the participants choose to mark the doctors gender in the no suggestions condition, they significantly more often mark the doctor as masc-coded than fem-coded with or without masc-coded suggestions (No suggestions: $t(144)=-3.862$, $p_{\text{FDR}}\approx 0.0012$; masc-coded suggestions: $t(316)=-8.542$, $p_{\text{FDR}}<0.0001$).
When it comes to fem-coded suggestions, we see significantly more fem-coded doctors with fem-coded suggestions than without any suggestions ($t(246)=3.219$, $p_{\text{FDR}}\approx 0.0081$). However, when shown fem-coded doctor suggestions, the difference between the rate of making the doctor character masc-coded vs fem-coded is not significantly different ($t(348)=-1.026$, $p_{\text{FDR}}\approx 0.5207$).
@@
Considering overall reliance, we find that participants type new words or edit a model-suggested word marginally significantly more in conditions where the model is prompted to make the doctor fem-coded (H2: $t(7692)=2.258$, $p_{\text{FDR}}\approx 0.0784$). When we only consider the story words that specify the doctor character’s gender (see example words in Table 19), we now see a significantly higher rate of overrides or edits in the fem-coded doctor conditions (H2a: $t(159)=2.734$, $p_{\text{FDR}}\approx 0.0294$). Similarly, when we consider only model suggestions that would specify the doctor character’s gender, we see a significantly higher rejection rate for words that would make the doctor fem-coded (H2b: $t(1014)=2.926$, $p_{\text{FDR}}\approx 0.0172$).
@@ A.1.2. Effects on Gender and Confidence
None $t(156)=-0.374$ $0.8283$
confident $t(156)=-2.857$ $0.0210$
M confident $t(156)=-0.139$ $0.9488$
@@
[…]
Breakdown of doctor gender and confidence in stories written with various suggestions. Confident suggestions of either gender pushes stories to have more gender-specified confident doctors. Unconfident masc-coded suggestions decrease the proportion of gender-unspecified confident doctors, but increases the proportion of confidence-unspecified masc-coded doctors not unconfident masc-coded doctors.
@@
We see that when the model provides unconfident masc-coded doctor suggestions, participants wrote significantly more masc-coded doctors of unspecified confidence ($t(155)=2.552$, $p_{\text{FDR}}\approx 0.0446$) and significantly fewer confident doctors of unspecified gender ($t(155)=-3.038$, $p_{\text{FDR}}\approx 0.0145$). No other set of attributes changed significantly in under these suggestions. One possible interpretation of these results is that participants take on the suggestion of the doctor being masc-coded, but largely refuse to make a masc-coded doctor unconfident and instead leave confidence unspecified.
We also see that when the model provides confident fem-coded doctor suggestions, participants wrote significantly more confident fem-coded doctors ($t(156)=2.91$, $p_{\text{FDR}}\approx 0.0195$) and significantly fewer confident doctors with gender unspecified, marginally more fem-coded doctors of unspecified confidence ($t(156)=2.194$, $p_{\text{FDR}}\approx 0.0955$), and significantly fewer confident doctors with unspecified gender ($t(156)=-2.857$, $p_{\text{FDR}}\approx 0.0210$). For confident masc-coded doctor suggestions, we see no significant changes, but note similar trends away from confident doctors of unspecified gender ($t(146)=-1.785$, $p_{\text{FDR}}\approx 0.1952$) and toward confident masc-coded doctors ($t(146)=1.975$, $p_{\text{FDR}}\approx 0.1402$). These results show that gendered confident suggestions are generally effective at shifting stories away from non-gendered confident doctors.
Overall, we note how regardless of suggestion type, we always continue to see a sizable group of stories about confident or unspecified masc-coded doctors, even when the opposite is suggested. On the other hand we see no fem-coded doctors when suggesting a confident masc-coded doctor and only see fem-coded doctors in a masc-coded doctor condition when the doctor is suggested to be unconfident. This again aligns with Cao et al. (2022b)’s findings about gender-confidence stereotypes.

### v2

@@ 6.4.2. Effects on Gender and Benevolence Jointly
Beyond gender on its own, we also consider the benevolence of the presidents (9(b)). For each configuration of suggestions and each potential set of attributes that could be given to the president character, we consider whether adding that suggestion type changes the proportion of presidents that have that set of attributes. We observe that overall, the model was not successful in convincing participants to make threatening president characters, with the rate of threatening president characters (regardless of gender) being quite low regardless of the presence or type of suggestions with $22/412$ stories containing a threatening president ($8/162$ for stories written with threatening suggestions of either gender).
We do see that when provided with benevolent or threatening fem-coded suggestions, participants wrote significantly more benevolent fem-coded presidents (benevolent fem-coded suggestions: $t(154)=3.068$, $p_{\text{FDR}}\approx 0.0134$, $d=0.492$; threatening fem-coded suggestions: $t(152)=2.526$, $p_{\text{FDR}}\approx 0.0465$, $d=0.407$). These results weakly support H1. Joint gender and benevolence suggestions generally did not change the distribution of the president characters’ attributes when comparing to stories written without suggestions. While fem-coded suggestions successfully increased the frequency of fem-coded president characters, participants did not accept these fem-coded characters being anti-stereotypically threatening, leading to an increase in benevolent fem-coded presidents when writing with either benevolent or threatening fem-coded suggestions when comparing to stories written with no suggestions.
@@
Bar charts showing character gender distributions under different suggestions for participants who rated straight women as more versus less competent than men. The distributions are similar across groups, with higher proportions of feminine-coded doctors and detectives under suggestions among participants who rated straight women as more competent.
6.5. Additional Effects: Time to Decide and Individual Differences
Beyond our primary hypotheses, here we consider the effects of suggestion type on the time to make individual decisions (expanded on in subsection B.1), effects of participant’s views and stereotypes on story attributes (expanded on with other individual differences in subsection B.3), and the effect of English proficiency on participants’ reliance on model suggestions. In the appendix, we further consider the effects of suggestions on overall story length and time to write (subsection B.2), the distribution and correlation of participant stereotypes (subsection B.4), and the effect of suggestions on toxicity, sentiment, and characters’ agency in the co-written stories (subsection B.5).
6.5.1. Effects of Suggestion Type on Time to Make Decisions
In the Detectives scenario, we find that participants take significantly less time to decide whether accept suggestions of the partner being masc-coded when the partner is untrustworthy as opposed to trustworthy (t(306)=3.544t(306)=-3.544, $p_{\text{FDR}}\approx 0.0030$, d=0.464d=-0.464). Further, when the partner is untrustworthy, it takes significantly less time for participants to make decisions about masc-coded suggestions than fem-coded ones (t(310)=3.48t(310)=-3.48, $p_{\text{FDR}}\approx 0.0036$, d=0.455d=-0.455). This could mean that masc-coded untrustworthy detective partners are the least unexpected group. This finding is in line with our findings about the rate of masc-coded untrustworthy detectives in stories written without suggestions and is also in line with cao-etal-2022-theory’s findings about gender-trustworthiness stereotypes.
@@ 6.5.1. Effects of Suggestion Type on Time to Make Decisions
     2 (Limited working) t(17403)=8.97t(17403)=-8.97 $0.0000$ *
     3 (Professional working) t(17673)=8.228t(17673)=-8.228 $0.0000$ *
     4 (Full professional) t(18053)=10.37t(18053)=-10.37 $0.0000$ *
     5 (Primary fluency/bilingual) t(25097)=8.451t(25097)=-8.451 $0.0000$ *
@@
     5 (Primary fluency/bilingual) $t(25568)=3.258$ $0.0066$ *
Figure 11. AI reliance broken down by self-reported English proficiency. (a) We break down the set of words in stories written by participants based on their source (e.g., the participant chose a model suggestion by pressing a suggestion button). (b) We compare the proportion of model suggested words to human written or edited words in pairs of proficiency levels.
Bar chart showing participants’ reliance on model suggestions across English proficiency levels. Participants in the “elementary proficiency” category show higher rates of accepting suggestions, either by pressing the suggestion button or by typing the model suggestion exactly.
@@ 6.5.2. Effect of Participants’ Views on Gender
Here we consider how participants’ views on gender affect the stories they write with and without suggestions (H4). Specifically, we focus on participants’ perceptions of the “competence” of people of different genders and how these perceptions influence their gendering of doctor, president, and detective characters. We hypothesize that participants who believe that women are more competent than men will write more stories about fem-coded doctors, presidents, and detectives without suggestions and be more likely to accept model suggestions of these characters being fem-coded.
We focus this analysis on the beliefs about straight men and women as participants likely defaulted to these characters being straight. We exclude the $42\%$ of participants who rated straight men and women’s competence within 10 points of each other (on a 0-100 scale) and then split participants into those who marked straight women as more competent ($53\%$ of the included participants) vs less competent ($47\%$ of the included participants) than straight men. We plot the breakdown of character gender for these two groups in each scenario in Figure 10.
For all scenarios, when we compare the gendered competence groups, we see no significant difference in the proportion of masc-coded or fem-coded characters written without suggestions or with masc-coded suggestions. We do, however, observe a higher rate of fem-coded characters when suggested in the Doctor and Detectives scenarios. This trend is significant for the Detectives scenario ($t(97)=2.742$, $p_{\text{FDR}}\approx 0.0304$, $d=0.555$) and marginally significant for the Doctor scenario ($t(103)=2.279$, $p_{\text{FDR}}\approx 0.0802$, $d=0.445$). This provides some evidence that participants are more willing to accept anti-stereotypical suggestions when they (or their close friends) hold anti-stereotypical beliefs.
As we discuss in more detail in subsection B.3 in the appendix, we do not find evidence that participants’ gender identity directly affects acceptance of gendered suggestions. However, we do find that men are more likely than women to endorse the belief that men are more competent than women ($t(412)=2.082$, $p\approx 0.0379$, d=0.209d=0.209). [^12] This test was exploratory, not part of the main analysis, and was not pre-registered. As such, it was excluded from the Benjamini-Hochberg correction applied to the primary analyses. These results highlight that binary gender identities do not capture a uniform set of experiences or beliefs. In contrast, measures of gender-related attitudes, though not uniformly predictive across writing scenarios, offer comparatively more insight into participants’ interactions with gendered predictive text suggestions.
@@ 7. Discussion, Limitations, and Implications
In this work, we examined the effect of biased predictive text suggestions on human-AI co-written text. Predictive text is widely used in mobile interfaces like the one examined in this study. These systems are not neutral. The underlying models may produce gender-biased suggestions that reflect or reinforce social stereotypes. When people accept biased suggestions, the resulting co-written texts may perpetuate these stereotypical associations, potentially shaping the beliefs of those who read them and those who wrote them. This is especially concerning for children who are still forming their beliefs about the world (Peterson and Lach, 1990) and for non-native speakers who generally accept more model suggestions (Buschek et al., 2021). These biases may also create feedback loops: if human-AI co-written texts containing gender stereotypes are later used to train future models, even an initially “unbiased” system could become increasingly biased over time, a challenge that would not be solved with watermarking because the resulting text is human-written.
Our findings show that people are not equally influenced by pro-stereotypical and anti-stereotypical suggestions. While anti-stereotypical suggestions can, in some contexts, increase the proportion of anti-stereotypical writing, this is often not consistent enough to offset pro-stereotypical human biases. This pattern contrasts with prior research showing that model suggestions can steer writing in multiple directions (e.g., positive vs negative sentiment, arguing social media is good vs bad for society, etc). The contrast between our findings and findings from prior work may be due to the stickiness of stereotypes that people hold, which are often less malleable than overt beliefs about the world (Dasgupta and Greenwald, 2001; Lai et al., 2014; Payne and Gawronski, 2010; Howe and Krosnick, 2017; Lewandowsky et al., 2012). It may also be that single-word suggestions, as used in our study, have less influence on co-writing outcomes than sentence- or paragraph-level suggestions (Dhillon et al., 2024).
Still, our work demonstrates that mitigating extrinsic bias in the model (or even producing only anti-stereotypical suggestions) may not lead to the sociotechnical mitigation of bias in the human-AI outcome. Even when developers create models that generate outputs aligned with certain fairness principles, human preferences and biases can reintroduce inequity through selective uptake of those suggestions. As a result, we should not expect co-writing with a “perfectly debiased” predictive text system to yield perfectly unbiased stories. For instance, interpolating between suggestion configurations, we estimate that even a model offering masc-coded and fem-coded suggestions at equal rates would yield only about $25.5\%$ fem-coded detectives. These differences in uptake echo prior research showing that users sometimes prefer gender-biased career recommendations that align with their own expectations, even when “debiased” alternatives are available (Krause et al., 2025; Wang et al., 2023a, 2022).
Our findings highlight that technical interventions to reduce model bias may not be sufficient to achieve equitable outcomes when humans and AI work together. The biases in the final co-written texts in our study come not solely from the model but from participants’ decisions to accept or override AI suggestions. For design, this suggests that fairness should be treated as a property of the human-AI system as a whole, not merely of the AI. One possible avenue for future HCI research is to explore designs that foster more engagement with anti-stereotypical suggestions, encouraging users to reflect on and occasionally challenge their own assumptions. Such systems could draw from existing work on implicit bias mitigation, which has developed techniques to help people recognize and reduce biased beliefs (though these mitigation methods may not have a long-term impact (Dasgupta and Greenwald, 2001; Lai et al., 2014; Payne and Gawronski, 2010)). By connecting technical fairness work with behavioral design strategies, future research could help bridge the gap between algorithmic debiasing and sociotechnical fairness in practice.
Our study has several limitations. Participants were asked to co-author a story that was not entirely their own. As a result, they may have lacked a clear narrative plan, potentially amplifying the influence of model suggestions. Some of the content in the provided story prefixes and controlled model suggestions were more indicative of US-centric cultural norms and biases (e.g., in the choice of character names), which made the tasks less realistic for participants not based in the United States. Moreover, while our study focused on a creative writing scenario, predictive text systems are often used in everyday communication, where the effects of bias may manifest differently. For writing that is more grounded in a real-world experience or interaction, a predictive text system may affect how an author describes a past appointment with a doctor (e.g., the doctor’s disposition or agency) but is unlikely to influence how one describes the doctor’s gender itself. The study further considered only single-word predictive text. While this interaction mode is common in mobile applications, the findings may not generalize to co-writing applications using longer suggestions. In addition, the study focused exclusively on writing in English and on gender and sexuality stereotypes that have been documented among people in the United States (cao-etal-2022-theory). Although our sample included some participants who were not native English speakers or based in the United States, their post-survey responses were broadly consistent with the belief patterns reported in prior U.S.-based work. Nonetheless, our results may not generalize to other cultural or linguistic contexts, or to stereotypes concerning other personal characteristics or sensitive attributes.
Overall, our work shows that anti-stereotypical predictive text suggestions have some potential to lessen gender and sexuality biases in human-AI co-writing, but these suggestions alone are not enough to encourage users to break out of stereotypical patterns. Pro-stereotypical narratives continue to dominate even under maximally anti-stereotypical system settings. We therefore caution against over-relying on purely technical debiasing as a fairness solution. Instead, we advocate for future HCI and AI design research that considers interventions at the interaction level, supporting users in reflecting on, engaging with, and potentially revising their own beliefs or biases during the act of co-writing. By attending to both model design and human behavior, we may better understand and shape the sociotechnical dynamics that produce bias in human-AI collaboration.
@@ A.1. Scenario: Doctor
In this scenario, the story prefix describes the speaker visiting the doctor. The model suggests that the doctor is of a particular gender and is either “confident” or “unconfident” (See examples in Table 5). According to cao-etal-2022-theory, American annotators view men as comparatively more confident than women, though our participants may not have held this same stereotype (See Table 16).
@@ A.1.1. Effects on Gender Alone
First, we consider how suggestions affect how participants specify a doctor character’s gender alone (Table 3). At the level of stories, we hypothesized that participants would default to making the doctor masc-coded, leading the no suggestions condition to be similar to the masc-coded suggestions condition. However, we do see marginally more masc-coded doctors when participants are given masc-coded suggestions than no suggestions ($t(230)=2.467$, $p_{\text{FDR}}\approx 0.0521$, $d=0.349$). We believe this may be due to an overall lower rate of specifying the doctor’s gender in the no suggestions condition. When the participants choose to mark the doctors gender in the no suggestions condition, they significantly more often mark the doctor as masc-coded than fem-coded with or without masc-coded suggestions (No suggestions: t(144)=3.862t(144)=-3.862, $p_{\text{FDR}}\approx 0.0012$, d=0.639d=-0.639; masc-coded suggestions: t(316)=8.542t(316)=-8.542, $p_{\text{FDR}}<0.0001$, d=0.958d=-0.958).
When it comes to fem-coded suggestions, we see significantly more fem-coded doctors with fem-coded suggestions than without any suggestions ($t(246)=3.219$, $p_{\text{FDR}}\approx 0.0081$, $d=0.449$). However, when shown fem-coded doctor suggestions, the difference between the rate of making the doctor character masc-coded vs fem-coded is not significantly different (t(348)=1.026t(348)=-1.026, $p_{\text{FDR}}\approx 0.5207$, d=0.11d=-0.11).
@@
Considering overall reliance, we find that participants type new words or edit a model-suggested word marginally significantly more in conditions where the model is prompted to make the doctor fem-coded (H2: $t(7692)=2.258$, $p_{\text{FDR}}\approx 0.0784$, $d=0.052$). When we only consider the story words that specify the doctor character’s gender (See example words in Table 19), we now see a significantly higher rate of overrides or edits in the fem-coded doctor conditions (H2a: $t(159)=2.734$, $p_{\text{FDR}}\approx 0.0294$, $d=0.432$). Similarly, when we consider only model suggestions that would specify the doctor character’s gender, we see a significantly higher rejection rate for words that would make the doctor fem-coded (H2b: $t(1014)=2.926$, $p_{\text{FDR}}\approx 0.0172$, $d=0.184$).
@@ A.1.2. Effects on Gender and Confidence
None t(156)=0.374t(156)=-0.374 $0.8283$
confident t(156)=2.857t(156)=-2.857 $0.0210$ *
M confident t(156)=0.139t(156)=-0.139 $0.9488$
@@
Left: Stacked bar charts showing doctor gender and confidence in stories under different suggestion conditions. Right: results of t-tests comparing the presence of these attributes with ad without suggestions. Confident suggestions of either gender are associated with higher proportions of gender-specified confident doctors. Unconfident masculine-coded suggestions correspond to lower proportions of gender-unspecified confident doctors and higher proportions of confidence-unspecified masculine-coded doctors.
@@
We see that when the model provides unconfident masc-coded doctor suggestions, participants wrote significantly more masc-coded doctors of unspecified confidence ($t(155)=2.552$, $p_{\text{FDR}}\approx 0.0446$, $d=0.408$) and significantly fewer confident doctors of unspecified gender (t(155)=3.038t(155)=-3.038, $p_{\text{FDR}}\approx 0.0145$, d=0.486d=-0.486). No other set of attributes changed significantly in under these suggestions. One possible interpretation of these results is that participants take on the suggestion of the doctor being masc-coded, but largely refuse to make a masc-coded doctor unconfident and instead leave confidence unspecified.
We also see that when the model provides confident fem-coded doctor suggestions, participants wrote significantly more confident fem-coded doctors ($t(156)=2.91$, $p_{\text{FDR}}\approx 0.0195$, $d=0.464$) and significantly fewer confident doctors with gender unspecified, marginally more fem-coded doctors of unspecified confidence ($t(156)=2.194$, $p_{\text{FDR}}\approx 0.0955$, $d=0.35$), and significantly fewer confident doctors with unspecified gender (t(156)=2.857t(156)=-2.857, $p_{\text{FDR}}\approx 0.0210$, d=0.456d=-0.456). For confident masc-coded doctor suggestions, we see no significant changes, but note similar trends away from confident doctors of unspecified gender (t(146)=1.785t(146)=-1.785, $p_{\text{FDR}}\approx 0.1952$, d=0.293d=-0.293) and toward confident masc-coded doctors ($t(146)=1.975$, $p_{\text{FDR}}\approx 0.1402$, $d=0.325$). These results show that gendered confident suggestions are generally effective at shifting stories away from non-gendered confident doctors.
Overall, we note how regardless of suggestion type, we always continue to see a sizable group of stories about confident or unspecified masc-coded doctors, even when the opposite is suggested. On the other hand we see no fem-coded doctors when suggesting a confident masc-coded doctor and only see fem-coded doctors in a masc-coded doctor condition when the doctor is suggested to be unconfident. This again aligns with cao-etal-2022-theory’s findings about gender-confidence stereotypes.